if a wheel that has 3000 revolutions while rolling in a straight line for 45000 in what is the radius of the wheel
Answer by venugopalramana(3286)

if a wheel that has 3000 revolutions while rolling in a straight line for 45000 in what is the radius( R SAY) of the wheel
PERIMETER OF WHEEL = 2*PI*R = DISTANCE TRAVELLED IN 1 REVOLN.
DISTANCE TRAVELLED IN 3000 REVOLNS.=3000*2PI*R = 45000
R = 45000/(3000*2*PI)=7.5/PI = 7.5/3.14= 2.4 APPROX